OpenStudy (anonymous):

Determine a region whose area is equal to the given limit. Do not evaluate the limit. lim┬(n→∞)⁡∑_(i=1 )^n 2/n (1+2i/n)^18

OpenStudy (zarkon):

\[\lim_{n\to\infty}\sum_{i=1}^{n}\frac{2}{n}\left(1+\frac{2i}{n}\right)^{18}\]

OpenStudy (zarkon):

\[\lim_{n\to\infty}\sum_{i=1}^{n}\frac{2}{n}\left(1+\frac{2i}{n}\right)^{18}\] \[\int\limits_{a}^{b}f(x)dx=\lim_{n\to\infty}\sum_{i=1}^{n}f\left(a+i\frac{b-a}{n}\right)\frac{b-a}{n}\]

OpenStudy (zarkon):

\[a=1\] \[b=3\] \[f(x)=x^{18}\]

OpenStudy (zarkon):

\[=\int\limits_{1}^{3}x^{18}dx\]

OpenStudy (zarkon):

so it is the area under the graph of \[f(x)=x^{18}\]between the x=1 and x=3

OpenStudy (anonymous):

Thats what I thought but webassign says that the answer is wrong

OpenStudy (zarkon):

the limit of the sum gives the same value as the integral

OpenStudy (zarkon):

\[\frac{1162261466}{9}\]

OpenStudy (zarkon):

over 19

OpenStudy (zarkon):

\[\frac{1162261466}{19}\]

OpenStudy (anonymous):

Here are the choices from webassign

OpenStudy (zarkon):

2nd one is equivalent to mine...do the substitution u=1+x

OpenStudy (anonymous):

I'm sorry, I'm not sure I understand

OpenStudy (zarkon):

\[\int\limits_{0}^{2}(1+x)^{18}dx\] let u=1+x du=dx then \[x=0\rightarrow u=1\] \[x=2\rightarrow u=3\] and the integral becomes \[\int\limits_{1}^{3}u^{18}du\] which is equivalent to what I wrote before